### v1

# A Synthetic Dataset for Manometry Recognition in Robotic Applications

Abstract
This work addresses the challenges of data scarcity and high acquisition costs for training robust object detection models in complex industrial environments, such as offshore oil platforms. The practical and economic barriers to collecting real-world data in these hazardous settings often hamper the development of autonomous inspection systems. To overcome this, in this work we propose and validate a hybrid data synthesis pipeline that combines procedural rendering with AI-driven video generation. Our methodology leverages BlenderProc to create photorealistic images with precise annotations and controlled domain randomization, and integrates NVIDIA’s Cosmos-Predict2 world-foundation model to synthesize physically plausible video sequences with temporal diversity, capturing rare viewpoints and adverse conditions. We demonstrate that a YOLO-based detection network trained on a composite dataset, blending real images with our synthetic data, achieves superior performance compared to models trained exclusively on real-world data. Notably, a 1:1 mixture of real and synthetic data yielded the highest accuracy, surpassing the real-only baseline. These findings highlight the viability of a synthetic-first approach as an efficient, cost-effective, and safe alternative for developing reliable perception systems in safety-critical and resource-constrained industrial applications.
I INTRODUCTION
The uninterrupted functioning and security of vital infrastructure, especially within the petroleum and natural gas sectors, are of utmost importance due to the inherent dangers associated with the handled materials and the severe repercussions of system malfunctions [1, 2, 3]. Conventional maintenance and inspection procedures in these intricate industrial settings are largely dependent on human personnel, who are frequently exposed to substantial hazards such as contact with toxic agents, extreme temperatures, and confined spaces [4, 5]. Accordingly, there is a strong demand for technologies that increase inspection effectiveness [6, 7] while simultaneously reducing human exposure in dangerous scenarios [8, 9, 10].
The introduction of autonomous legged robots, e.g., Unitree B2 [11] and ANYmal C [12], offers a promising alternative [13, 14]. Unlike wheeled or tracked systems that struggle with uneven terrain, legged platforms can operate in a reliable manner even in the presence of stairs and irregular surfaces [15], enabling routine, automated data capture that minimizes downtime and reduces human risk [16].
However, reliable perception remains a key bottleneck: robots must detect and localize facility-critical assets—valves, pipes, gauges, and safety equipment—to assess conditions and comply with procedures in visually cluttered plants [17, 3]. Acquiring and annotating large, diverse datasets for training modern detectors (e.g., YOLO) is time-consuming, costly, and often disruptive to operations; access constraints and safety risks further limit in-plant data collection [18, 19, 20, 21].
To overcome these limitations, in this work, we leverage high-quality synthetic data at two complementary levels. First, we use BlenderProc—an open-source, procedural rendering pipeline—to generate photorealistic images with controlled domain randomization over geometry, illumination, materials, and sensors [22, 23, 24]. Second, we integrate NVIDIA Cosmos-Predict2, NVIDIA’s latest world-foundation model (WFM) family for “physical AI” [25, 26, 27], via ComfyUI [28] workflows to synthesize physically plausible imagery and video priors from rare viewpoints, adverse lighting, and hard-to-stage events, thus broadening coverage beyond what is practical to render or capture in the field. Cosmos provides model variants and deployment tooling intended for robotics and AV pipelines, enabling scalable generation and curation that complements procedural graphics and reduces the sim-to-real gap in downstream perception [29].
This work introduces a practical pipeline that blends BlenderProc rendering with Cosmos-Predict2 generation to mitigate data scarcity for vision-based inspection in oil and gas facilities. We build a large-scale, richly annotated synthetic dataset and train a YOLO-based detector, showing that synthetic data can substantially reduce reliance on costly and hazardous in-plant collection while improving recognition of safety-critical assets in complex environments.
The rest of this paper is organized as follows. Section II presents the related works. Section III details the unified pipeline—procedural dataset creation with BlenderProc and AI-extended synthesis with Cosmos-Predict2 (ComfyUI)—and training/validation of the object detector. Section IV presents quantitative and qualitative results, and Section V concludes with limitations and future work.
II Related Work
Synthetic data has become a practical solution for the scarcity and safety constraints of in-plant data collection, with domain randomization and photorealistic rendering shown to reduce the sim-to-real gap for perception and control [30]. Early sim-to-real studies demonstrated that heavy appearance randomization can transfer object localization and even collision-avoidance policies trained purely in simulation to the real world [31, 32]. Recent analyses formalize when and how randomization narrows the gap and how to tune distributions over simulator parameters [33, 34].
On the procedural/graphics side, toolchains such as BlenderProc provide modular, scriptable generation of photorealistic images with rich annotations and sensor simulation—now a common backbone for dataset bootstrapping in robotics and industrial vision [22]. Parallel efforts in commercial-grade simulators (NVIDIA Omniverse/Isaac Sim Replicator) expose APIs for large-scale synthetic data generation with domain randomization, sensor models, and annotators, frequently used in AV/robotics pipelines [35, 36]. Unity’s Perception stack similarly targets turnkey dataset creation with labelers and randomized scenes for object detection and pose estimation [37].
Complementary to procedural rendering, world foundation models (WFMs) have emerged as AI-native generators that simulate and predict physically plausible scenes directly from text/visual conditions. NVIDIA’s Cosmos platform targets “physical AI” use cases (robotics, AV, industrial vision), exposing Cosmos-Predict2 models for Video-to-World generation (2B and 14B variants) and post-training/fine-tuning workflows aimed at scalable coverage of rare viewpoints, adverse illumination, and hard-to-stage events [38, 39, 40]. Public releases and reports emphasize open accessibility and training on large corpora of robotics/drive videos, positioning Cosmos as a data engine to complement graphics-based pipelines rather than replace them [41, 42]. Community workflows via ComfyUI further lower integration friction, enabling mixed procedural+generative curation where Cosmos assets are inserted into Blender/YOLO pipelines [43]. Taken together, recent work indicates that hybrid pipelines, procedural rendering for precise labels and controllable physics, plus WFM-based synthesis for broader coverage and temporal variation, can materially reduce reliance on hazardous, costly field acquisition in complex industrial environments.
III Methodology
Our methodology combines real data acquisition, synthetic data generation through BlenderProc, and AI-extended video synthesis using NVIDIA Cosmos Predict2 integrated via ComfyUI. The unified pipeline is illustrated in Fig. 1.
Figure 1: Proposed methodology pipeline for dataset creation and training. The framework integrates three sources: (i) real data acquisition and annotation (manual via CVAT and semi-automatic via SAM2), (ii) synthetic renders generated in BlenderProc with domain randomization (backgrounds, illumination, pose, post-processing), and (iii) AI-extended video frames generated through ComfyUI with NVIDIA Cosmos Predict2 (relighting, motion, occlusion). All data sources are reviewed, merged, and used for training the YOLO-based object detector.
III-A Real Data Acquisition and Annotation
We collected approximately 2,500 real images and video frames of industrial manometers. These samples were annotated using a hybrid approach: manual bounding-box labeling with CVAT and semi-automatic segmentation with SAM2, as shown in Fig. 2. A small portion of these annotations was further reviewed by experts to ensure quality.
(a) Manual bounding boxes in CVAT.
(b) SAM2-assisted segmentation overlays.
Figure 2: Examples of real data annotations: (a) manual boxes curated in CVAT; (b) semi-automatic pixel masks using SAM2, later spot-checked by experts.
III-B Synthetic Data with BlenderProc
3D models of manometers were rendered with BlenderProc. To improve generalization and reduce overfitting, domain randomization was applied: random backgrounds (industrial textures and photos), lighting variations, camera pose sampling, and post-processing (noise, blur, chromatic aberration), with examples provided in Fig. 3. BlenderProc provides pixel-perfect labels, ensuring accurate segmentation and bounding-box masks.
(a) Manometter in a random background.
(b) An example of multi .blend files.
Figure 3: BlenderProc synthetic samples.
III-C AI-Extended Video Generation (ComfyUI + Cosmos Predict2)
To complement static synthetic renders, we integrated AI-driven video extension, illustrated in Fig. 4. Real short clips of manometers were expanded using ComfyUI workflows with NVIDIA Cosmos Predict2, enabling frame synthesis with temporal consistency, relighting, viewpoint changes, and motion blur. These pseudo-extended frames introduce temporal diversity (occlusion, reflections, jitter) that is difficult to model with static rendering. Pseudo-labels were propagated across frames using tracking and filtered with confidence thresholds, with human-in-the-loop auditing on a subset to mitigate noise.
(a) Relighting and viewpoint changes with motion blur.
(b) A frame from a shot clip made with Cosmos.
Figure 4: AI-extended video frames produced via ComfyUI + NVIDIA Cosmos Predict2.
III-D Data Composition Strategy
All data sources were merged into a unified dataset. To investigate the impact of proportions, we defined scenarios with:
  • Real only (2,500 images)
  • Mixed (real + synthetic in ratios 1:1, 1:3)
Within the synthetic portion, we tested different splits between BlenderProc renders and AI-extended video frames, adopting 70% BlenderProc and 30% video-based synthesis as baseline.
III-E Training and Evaluation
All scenarios were trained with the same YOLO-based architecture using identical hyperparameters. Evaluation was performed exclusively on a held-out set of real images. Metrics include mAP@[.5:.95], recall, average recall (AR), per-class AP, and PR curves. We also report normalized metrics (mAP divided by average instances per image), shown in Fig. 7, and perform ablations on domain randomization factors.
IV RESULTS
In this section we present the experimental protocol, justification for synthetic data generation, proportionality choices between real and synthetic samples, and the results obtained in different scenarios.
IV-A Experimental Protocol
Four training scenarios were defined: (i) only real images (2,500) and (ii) a mix of real and synthetic images.
For mixed datasets, we tested ratios of 0.5:0.5, 1:1, and 1:3 (real:synthetic). All scenarios were trained with the same YOLO-based detection architecture and evaluated on the same test split of real images.
IV-B Justification
The adoption of synthetic data is motivated by three factors: 1) the cost and time of annotating real images are significantly higher than generating synthetic ones; 2) domain randomization techniques such as random backgrounds, lighting and camera poses reduce overfitting and bridge the sim-to-real gap; 3) prior work shows that mixing 5–20% real data with synthetic data is often enough to reach competitive performance.
IV-C Proportionality of Synthetic Data
Considering the availability of 2,500 real photos, we defined proportionality as follows: - 1:1 → 2,500 real + 2,500 synthetic - 1:3 → 2,000 real + 7,500 synthetic
Additionally, a synthetic-only set was created with 20,000 images to serve as control set.
IV-D Results
Table I shows dataset composition and Table II reports object detection performance. Fig. 5.
TABLE I: Dataset composition (examples).
Scenario Real Images Synthetic Images Total
Real only 2500 0 2500
Mix 1:1 2500 2500 5000
Mix 1:3 2500 7500 10000
Mix 0.5:0.5 1250 1250 2500
TABLE II: Performance comparison (updated with 0.5:0.5).
Scenario mAP@[.5:.95] Recall AR
Real only 0.936 0.969 0.969
Mix 1:1 0.962 0.972 0.972
Mix 1:3 0.928 0.948 0.948
Mix 0.5:0.5 0.868 0.933 0.933
These results indicate that mixed datasets outperform the real-only baseline. In our experiments, the 1:1 mix achieved the highest mAP50-95, while the 1:3 mix still surpassed the real-only model but trailed the 1:1 configuration, as depicted in the training evolution in Fig. 6. These findings corroborate the broader consensus that blending a modest amount of real data with a larger volume of synthetic data can guide the detector toward realistic features while benefiting from the diversity and volume of synthetic samples.
Figure 5: Object instances per dataset scenario across four settings (Real only, Mix 1:1, Mix 1:3, Mix 0.5:0.5). Synthetic augmentation increases instance density; notably, Mix 0.5:0.5 uses only half the real images yet preserves a high number of instances, supporting robust training despite reduced real data.
Figure 6: mAP@[.5:.95] vs epochs for all scenarios. The Mix 1:1 configuration consistently achieves the best accuracy. The Real-only and Mix 1:3 curves converge to very similar performance, while Mix 0.5:0.5 underperforms. These results indicate that adding synthetic data is beneficial, but excessive synthetic proportion (e.g., 1:3) does not yield further gains over Real-only and incurs additional compute cost.
Figure 7: Normalized mAP (0.5–0.95) per dataset scenario, corrected by the average number of instances per image.
IV-E Additional Analyses
Figure 8: Precision over training epochs for four training regimes. The synthetic-augmented configuration reaches high precision within the first epochs and remains near saturation. Synthetic data can accelerate the reduction of false positives.
Figure 9: Recall across training epochs for four training regimes. Mix 1:1 provides the most consistent recall gains. Real-only and Mix 1:3 display similar recall trajectories, whereas Mix 0.5:0.5 is lower yet notably better than a naive halving of the real dataset would suggest, evidencing the compensatory effect of synthetic data.
Figure 10: F1-score over epochs for four training regimes. The curves mirror the recall advantage of Mix 1:1 while preserving high precision. Real-only and Mix 1:3 are nearly tied, and Mix 0.5:0.5 trails but benefits from synthetic augmentation given the reduced real data.
Figure 11: Early-epoch mAP50(B) (first 20 epochs) comparing four training regimes.
Figures 810 showcase the training dynamics of Precision, Recall, and F1-score across the four regimes (Real only, Mix 1:1, Mix 1:3, and Mix 0.5:0.5), enabling a side-by-side inspection of convergence behavior. A consistent pattern is that the Mix 1:1 configuration accelerates early learning relative to the Real-only baseline, reflected in higher early-epoch averages (e.g., absolute gains of $\sim$+0.024 in Precision, $\sim$+0.042 in Recall, and $\sim$+0.034 in F1 within the first 20 epochs). In contrast, Mix 1:3 tends to underperform in the same interval and later converges to results that are practically tied with Real only. The Mix 0.5:0.5 setting exhibits the slowest convergence and lowest asymptotic accuracy; nevertheless, despite using only half the real images, synthetic augmentation prevents a sharp degradation, indicating that synthetic data can partially compensate for small or imperfect real datasets. Complementarily, Fig. 11 highlights the early-epoch window of mAP50(B) (up to epoch 20): Mix 1:1 surpasses the Real-only model (by $\sim$+0.03 absolute mAP50 at the end of the window), Mix 1:3 hovers near or slightly below the baseline, and Mix 0.5:0.5 remains lower throughout. Taken together, these analyses suggest that synthetic data can materially improve early optimization and stability, especially in Recall and the resulting F1, provided that the synthesis pipeline is properly configured; otherwise, suboptimal choices in synthetic generation may negate the expected benefits.
IV-F Synthetic Data Composition: BlenderProc vs. Video-based
Inside the synthetic portion, we explored the impact of splitting between BlenderProc-rendered images and AI-extended video frames. BlenderProc provides proper labels and strong control over background, illumination, and pose, while video-based synthesis injects temporal variation such as blur, occlusion, and camera jitter. The baseline adopted was $70\%$ BlenderProc and $30\%$ video-based frames, which balances labeling accuracy and temporal diversity.
V CONCLUSIONS
This study demonstrates that mixing synthetic data with real images improves detection performance and training dynamics in industrial scenarios. Among the tested configurations, the 1:1 mixture delivers the best overall accuracy. Real-only and 1:3 converge to practically the same performance, indicating diminishing returns when the synthetic portion overwhelms the real data. The 0.5:0.5 setting underperforms the other mixes, but—despite using only half the real images—synthetic augmentation prevents a sharp degradation, suggesting that synthetic data can effectively compensate for small or imperfect real datasets.
Future work will extend the framework to more complex scenes, incorporate multimodal data (e.g., depth and infrared), and evaluate cross-domain transferability. This methodology has the potential to contribute to advancing robust perception systems with lower development costs and broader applicability.
Key takeaways: (i) The best-performing network is the 1:1 mix (real + synthetic); (ii) Real-only and 1:3 are practically tied at convergence, showing limited gains from excessive synthetic proportion; (iii) The 0.5:0.5 mix underperforms but remains viable—synthetic data mitigates the drop expected from halving real images; (iv) Beyond accuracy, heavier synthetic mixes increase compute cost without commensurate gains, so moderate synthetic ratios are preferable for cost–benefit; (v) Synthetic data is particularly valuable to “patch” small or faulty real datasets, stabilizing training and preserving generalization.
VI DATA AVAILABILITY
The curated dataset supporting this study (images, annotations, and train/validation/test splits) is publicly available on Figshare at Synthetic Data Manometry Interpretation  [Dataset]. [44]. Version 1.0, released Aug. 2025, licensed under CC BY 4.0.